Susan had $20,000 to invest. She decided to invest part of it in a high-risk bond that yields 8.2% annual interest, and the rest in a savings bank yielding 3.9% per year. What least amount should she invest in the high-risk bond if she wants to receive at least $1,000 interest per year?
Answer by mananth(16946)

Investment II 8.20% per annum ---x
Investment II 3.90% per annum ---y

x + y= 20000 ------------------------1
8.20% x + 3.90% y= = $1,000.00
Multiply by 100
8.2 x + 3.9 y= = $100,000.00 --------2

Multiply (1) by -8.2
we get

-8.2 x -8.2 y= = -164000.00

Add this to (2)

0 x -4.3 y= = -$64,000.00

divide by -4.3

y = $14,883.72 investment at
Balance $5,116.28 investment at
CHECK
$5,116.28 --------- 8.20% ------- $419.53
$14,883.72 ------- 3.90% ------- $580.47
Total ------- $1,000.00